OpenStudy (anonymous):

Identify the equation and graph of the line passing through point (2, -4) and having a slope of 3/2

OpenStudy (anonymous):

3x-2y=14 3x+2y=-2 -2x+3y=-16 2x+3y=-8

OpenStudy (anonymous):

So you have a point, and the slope, just graph it! find out other points and plot it :)

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

What's the slope :)

OpenStudy (anonymous):

m=3

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

I will be right back :)

OpenStudy (anonymous):

so x is a 3?

OpenStudy (anonymous):

M= 3/2

OpenStudy (anonymous):

In need a graph

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Plug them into Point-Slope form. \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value of the point, \(x_1\) is the x-value of the point, and \(m\) is the slope. So we have: \(y + 4 = \dfrac{3}{2}(x - 2)\) Distribute 3/2 into the parenthesis: \(y + 4 = \dfrac{3}{2}x - 3\) Subtract 4 to both sides, what's -3 - 4?

OpenStudy (anonymous):

I think its 3x-2y=14

OpenStudy (anonymous):

-7

OpenStudy (igreen):

Yup, you're correct.

OpenStudy (igreen):

3x - 2y = 14 is correct.

OpenStudy (anonymous):

cool

OpenStudy (igreen):

Here's a graph: https://www.desmos.com/calculator/r2tfxbrlco